OpenStudy (anonymous):

I need help A certain number is added to 6 and the result is multiplied by 25. The final answer is 50. Find the number?

OpenStudy (anonymous):

\[\left( x+6 \right)*25=50\] find x

OpenStudy (a1234):

Set up an equation as (x + 6) * 25 = 50. Use the distributive property.

OpenStudy (anonymous):

\[x+6=\frac{ 50 }{ 25}=2,x=?\]